OpenStudy (anonymous):

2 log x- 3 log (y+1) write the given expression as a single logarithm

OpenStudy (anonymous):

two steps for this one \[n\log(x)=\log(x^n)\] and \[\log(x)-\log(y)=\log(\frac{x}{y})\]

OpenStudy (anonymous):

use this \[\log a - \log b = \log a/b\] \[a \log b = \log b ^{a}\]

OpenStudy (anonymous):

how do u work it out?

OpenStudy (anonymous):

\[2\log x = \log x ^{2}\] for example

OpenStudy (anonymous):

oops \[3\log(y+1)=\log((y+1)^3)\]

OpenStudy (anonymous):

dont u also get log ^x2/ (y+1)^3

OpenStudy (anonymous):

???